One side of a rectangle is 7 inches shorter than 2 times the other side find the dimensions if the perimeter is 64
Answer by Cromlix(4381)

Hi there,
Make other side = x
One side = 2x - 7
Perimeter of a rectangle
= 2 * x + 2 * 2x - 7 (* times)
64 = 2x + 4x - 14
Collect like terms
2x + 4x = 64 + 14
6x = 78
x = 13
Other side = 13 inches
One side = 2(13) - 7
One side = 26 - 7
One side = 19
One side = 19 inches
Hope this helps :-)